OpenStudy (anonymous):

Brian has scored 79, 87, 61, 89, and 79on his previous five tests. What score does he need on his next test so that his average (mean) is 80?

OpenStudy (kira_yamato):

If \[\mu = 80\]\[\Sigma x = 80*6 = 480\]\[480 = 79+87+61+89+79+x \rightarrow x = 85\]

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

From where you got 6..??

OpenStudy (kira_yamato):

Because there is a total of 6 tests for the mean to be 80. Therefore n = 6, and ∑x = nµ

OpenStudy (anonymous):

Ok